Journalist Report – February 19th

Sol 7 – Winter is over, time for a Spring cleanup!

« “We won’t ruin Mars,” said the captain. “It’s too big and too good.”

“You think not? We Earth Men have a talent for ruining big, beautiful things.” »

– Chapter 7 of The Martian Chronicles by Ray Bradbury

After dancing, laughing, and singing our hearts out and playing card games over hot chocolate on the night of Sol 6, we allowed ourselves to sleep until 8:30 on this beautiful Sunday morning. Corentin had prepared an intense workout session to drain us completely and give us a fresh start for the incoming week. We had a lot of fun pushing ourselves close to and beyond our limits for activities such as push-ups, squats and pull-ups; it had been a while since our last truly exhausting and body-resetting workout.

Sunday at MDRS also rhymes with cleaning: after eating freshly baked pancakes for breakfast, we were set on giving the Hab a real makeover. After two very technical EVAs, the Lower Deck was in need of tidying, and the Upper Deck’s kitchen top gave us the opportunity to complete the second workout of the day. We were very proud to have used only 4 liters of water to do the dishes today, given the number of pots and pans left uncleaned after last night’s dinner and the pancakes! Since the beginning of the mission, we have managed to keep our water consumption relatively low, using about 8 liters per day per crewmember (compared to about 150 for the average American!). By 3pm, the Hab looked like it was brand new. Afterwards, Quentin and Alexandre played chess, Alice set her mind to making a new jump rope for our morning workouts, and Corentin and Jérémy installed two flags in the Hab: The United Nations flag and the flag of Humanity. Next to the latter, we placed the whiteboard on which we’ve been writing down quotes for the entire crew to read since the beginning of the mission. Today’s quote read: “Dreams are not what you have when you sleep. The true dreams are the ones that don’t let you sleep.” Food for thought!

In the unusually quiet Upper Deck of the Hab, with only a few muffled voices and music coming from our rooms, Alexandre stood looking out of the window and voiced how beautiful Mars was at this hour: a light pink hue tainted the mountains, reverberating on the last lingering patches of snow. It reminded me of what I have temporarily left behind to join in on this month-long mission: a damaged, mistreated, and disrespected Earth. Let this mission and Bradbury’s quote be a reminder to me and to all of us that protecting and healing our planet must be our top priority!

GreenHab Report – February 19th

GreenHab Officer: Adrien Tison

Environmental control: Heater

Average Temperature: 96.5 °F (at 12:00) & 77.9 °F (in the afternoon)

Average humidity: 14 % (to be double-checked)

Hours of supplemental light: 4 hours

Daily water usage for crops: 13.74 gallons

Daily water usage for research and/or other purposes: 0 gallons

Water in the blue tank: 108.23 gallons

Time(s) of watering for crops: 12:34 & 16:55

Changes to crops: N/A

Narrative: Yesterday, at 23:00 I noticed that the supplemental light wasn’t on, so I went to check on it and it seems like the timer was late. I then put the timer back on track.

When I entered the greenhab today, I was struck by how hot it was, indeed it was 96°F. Hopefully plants and crops were not too dry, but it was blatant they needed more water than usual.

To drop the temperature, I decided to attach the sunshade 50% on the right side and 30% on the left side. This way, the sun didn’t hit as much, and the temperature decreased.

After all of that, I maintained and watered crops and plants as my daily operations and got to harvest a whole lot of tomatoes!

Harvest: 305g of tomatoes

Support/supplies needed: N/A

Operations Report – February 19th

Crew 275 Operations Report 19Feb2023
SOL: 7

Name of person filing report: Quentin Royer
Non-nominal systems:
• Toilet
• 1 radio from the GreenHab

Notes on non-nominal systems: None

Spirit rover used: No
• Hours: 209.0
• Beginning charge: 100%
• Ending charge: N/A
• Currently charging: Yes

Opportunity rover used: No
• Hours: 113.1
• Beginning charge: 100%
• Ending charge: N/A
• Currently charging: Yes

Curiosity rover used: No
• Hours: 220.9
• Beginning charge: 100%
• Ending charge: N/A
• Currently charging: Yes

Perseverance rover used: No
• Hours: 255.5
• Beginning charge: 100%
• Ending charge: N/A
• Currently charging: Yes

General notes on rovers: Rovers were not used today, as there was no EVA.

Summary of Hab operations: The Crew tidied up the entire Hab. The Crew cleaned the Hab. 3 to 4 drops of water fell from the ceiling of the Upper Deck during the day – origin unknown.
• WATER USE: 18.02 gallons
• Water: 365.65 gallons
• Static tank pipe heater: On
• Static tank heater: On
• Toilet tank emptied: Yes

Summary of internet: Hughes Net was used during the day before the opening of COMMS for email purposes, and Starlink (MDRSCampus-Guest) is used during COMMS.

Summary of suits and radios: Nothing to report.

Summary of GreenHab operations: The Crew botanist conducted routine operations and harvested some tomatoes. He decided to use the sunshades in the afternoon in order to reduce the high temperature in the GreenHab
• WATER USE: 13.74 gallons
• Heater: On
• Supplemental light: From 11:00 pm to 2:00 am – The Crew botanist had to manually setup the timer at 11:00 pm after noticing that the supplemental light was off.
• Harvest: 305 g of tomatoes

Summary of ScienceDome operations: The Crew Scientist cleaned the ground and tidied up the equipment used during the week.
• Dual split: Switched off during the night

Summary of RAM operations: No operations were conducted in the RAM.

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: A Crew Member slightly hurt their back’s skin against a wall during this morning’s sport session. This was handled by HSO and there is absolutely nothing to worry about.

Questions, concerns, and requests to Mission Support:
• Did you troubleshoot the GreenHab radio? If so, is it in the RAM? We didn’t have any opportunity to go to the RAM today to check if it was in there.
• As the toilet’s tank replacement was postponed to the end of our rotation (thank you very much for that), should I remove “toilet” from the non-nominal systems or should I keep it?

Operations Report – February 19th

Crew 275 Sol Summary Report 19Feb2023

Sol: 7

Summary Title: Winter is over, time for a Spring cleanup!

Author’s name: Jeremy Rabineau, Commander.

Mission Status: Nominal.

Sol Activity Summary:

Yesterday evening, we took some time to celebrate the end of this successful first week at the MDRS. This was a great bonding time with games, music, and even some dancing! In the morning, we all took part in a longer sport session and then enjoyed nice pancakes together. We could tell by the crewmembers’ faces that everyone was happy to take some time away from the stress of the operations and to talk about other things. We also took this opportunity to fully clean and tidy up the station. Only our GreenHab Officer could not take the opportunity to change his schedule too much: no Sundays for the plants! Our batteries are now fully charged to start Week 2 with a lot of energy (and we will need it!).

Look Ahead Plan:

Tomorrow, we would like to start the first photogrammetry EVA at North Ridge. The aim will be to use drones to take pictures of the area and to postprocess them using photogrammetry. Then we will have a 3D map that will allow us to: 1) Contextualize the collection of rock samples; 2) Evaluate the added value of such maps to ease the exploration during the next EVA in this same area. On our way to North Ridge, we also plan to do some final adjustments, change the batteries, and retrieve the data of the atmospheric instruments located close to Marble Ritual.
After interacting with the researchers responsible of the MegaAres instrument, we came up with a plan for the maintenance of the MegaAres antenna. It will require a day with two EVAs next week: a short EVA for the retrieval of the antenna in the morning, and a longer one to put it back in place and perform some tests in the afternoon.

Anomalies in work: None.

Weather: Sunny, low -9°C/15°F, high 4°C/39°F.

Crew Physical Status: Optimal.

EVA: None.

Reports to be filed: Sol Summary, EVA Request, Operations Report, Journalist Report, GreenHab Report, HSO Report.

Support Requested: None.

Copyright © The Mars Society. All rights reserved. | Main Site